OpenStudy (21younga):

what is 2 5/7b - 2/3 A. 2 1/7 B. 2 1/21 C. 2 3/4 D. 1 3/4

OpenStudy (21younga):

PLEASE HELP I'VE USED A CALCULATOR AND IT STILL DOESN'T GET THE ANSWER RIGHT!!

OpenStudy (21younga):

I will medal for correct answer!

OpenStudy (whpalmer4):

Is that supposed to be \[2 \frac{5}{7} - \frac{2}{3}\]

OpenStudy (21younga):

yeah

OpenStudy (whpalmer4):

okay, when you used a calculator, what result did you get?

OpenStudy (21younga):

16/21

OpenStudy (whpalmer4):

okay, first convert \[2\frac{5}{7}\]to an improper fraction, what is the result?

OpenStudy (whpalmer4):

if you divide 1 into 7 equal parts, how many 1/7's do you have?

OpenStudy (whpalmer4):

if you divide a candy bar into 7 equal parts, you have 10 little pieces of candy bar?!?

OpenStudy (whpalmer4):

don't you have 7?

OpenStudy (21younga):

IDK

OpenStudy (whpalmer4):

that's okay, we don't need the computer, just your brain. you have an apple. you divide it into 4 equal parts. isn't each part 1/4 of the apple?

OpenStudy (21younga):

lol I read your response wrong

OpenStudy (whpalmer4):

okay, so if we convert 1 into 7ths, don't we have \[1 = \frac{7}{7}\]?

OpenStudy (21younga):

yes

OpenStudy (whpalmer4):

and then if we add another \(1\), it would be \[1+1 = \frac{7}{7} + \frac{7}{7} = \frac{14}{7}\] and then we add another \(\dfrac{5}{7}\) \[2\frac{5}{7} = 1+1+\frac{5}{7} = \frac{7}{7}+\frac{7}{7}+\frac{5}{7} = \frac{19}{7}\] agree, disagree?

OpenStudy (21younga):

agree

OpenStudy (whpalmer4):

in general, if you have a mixed fraction of the form \[a\frac{b}{c}\]you convert it to an improper fraction by doing \[a\frac{b}{c} = \frac{a*c + b}{c}\] So\[2\frac{5}7 = \frac{2*7+5}{7} = \frac{19}{7}\]

OpenStudy (whpalmer4):

so now our problem looks like \[\frac{19}7 -\frac{2}3\]how do you proceed from here?

OpenStudy (21younga):

you find the common denominator

OpenStudy (whpalmer4):

excellent. what is the common denominator you are going to use here?

OpenStudy (21younga):

21

OpenStudy (whpalmer4):

now we're talking. what is the problem going to look like after you convert both fractions to use 21 as the denominator?

OpenStudy (21younga):

57/21 -14/21

OpenStudy (whpalmer4):

good, so the final answer is....drumroll please...

OpenStudy (21younga):

43/21 or 2/21?

OpenStudy (21younga):

no 2 1/21

OpenStudy (21younga):

that's what i meant

OpenStudy (whpalmer4):

got it!

OpenStudy (whpalmer4):

any questions about how we did that?